# Uncovering Safety Risks of Large Language Models through Concept Activation Vector

### v3

@@ Abstract
Despite careful safety alignment, current large language models (LLMs) remain vulnerable to various attacks. To further unveil the safety risks of LLMs, we introduce a Safety Concept Activation Vector (SCAV) framework, which effectively guides the attacks by accurately interpreting LLMs’ safety mechanisms. We then develop an SCAV-guided attack method that can generate both attack prompts and embedding-level attacks with automatically selected perturbation hyperparameters. Both automatic and human evaluations demonstrate that our attack method significantly improves the attack success rate and response quality while requiring less training data. Additionally, we find that our generated attack prompts may be transferable to GPT-4, and the embedding-level attacks may also be transferred to other white-box LLMs whose parameters are known. Our experiments further uncover the safety risks present in current LLMs. For example, we find that six out of seven open-source LLMs that we attack consistently provide relevant answers to more than 85% malicious instructions. Finally, we provide insights into the safety mechanism of LLMs.
@@ 1 Introduction
The rapid advancement in large language models (LLMs) has raised significant concerns about their potential misuse  (Wei et al., 2024; Wang et al., 2024; Kang et al., 2023; Yao et al., 2024). Developers usually conduct intensive safety alignment works (Bai et al., 2022; Ouyang et al., 2022; OpenAI, 2024; Ge et al., 2023) to prevent powerful LLMs from being exploited for harmful activities. However, researchers have discovered that these time-consuming safety alignment efforts can be reversed by various attack methods (Liu et al., 2023; Zou et al., 2023b; Shen et al., 2023; Zou et al., 2023a). These methods can identify vulnerabilities in safety alignment technologies and enable developers to fix them promptly, reducing the societal safety risks of LLMs.
@@
Existing attack methods utilize different levels of information from LLMs to achieve varying degrees of model understanding and control. Pioneering attack methods manually design prompt templates (Shen et al., 2023; Li et al., 2023) or learn attack prompts without information about intermediate layers of LLMs (Liu et al., 2023; Zou et al., 2023b). The attack prompts may be applied to various LLMs, supporting both black-box attacks to APIs and white-box scenarios where model parameters are released. However, their attack success rates (ASR) (Zou et al., 2023b) are constrained by an insufficient understanding of LLMs’ internal working mechanisms. Some recent attack works further utilize model embeddings at intermediate layers (Zou et al., 2023a; Li et al., 2024). By better understanding models’ safety mechanisms and perturbing relevant dimensions in the embeddings, these methods achieve significantly higher ASR on white-box LLMs. However, they cannot be applied to black-box APIs. Moreover, existing methods perturb LLM embeddings based on potentially misleading heuristics (Section 2.3.1). Due to the lack of a principled optimization goal, they result in a suboptimal ASR, may generate low-quality (e.g., repetitive) text, and require time-consuming grid search to find a good combination of hyperparameters (e.g., perturbation magnitude and layers).
@@
We aim to solve the aforementioned issues by more accurately modeling LLMs’ safe mechanisms, based on which principled optimization goals can be developed to well balance ASR and response quality, enable automatic hyperparameter selection, and support both embedding- and prompt-level attacks. Specifically, we make the following contributions.
@@
First, we establish a Safety Concept Activation Vector (SCAV) framework that effectively guides the subsequent attack process by accurately interpreting LLMs’ safety mechanisms. It quantifies the probability that an LLM considers an embedding as malicious based on the Concept Activation Vector (Kim et al., 2018), which linearly separates embeddings of malicious and safe instructions. We show our safety mechanism modeling eliminates potential misleading heuristics of the attack baselines.
@@
We then develop an SCAV-guided attack method, which utilizes SCAV to design principled optimization goals for finding high-quality embedding- and prompt-level attacks. Our embedding-level attack method eliminates the time-consuming hyperparameter tuning process by providing a closed-form solution of the optimal perturbation magnitudes and automatically selecting layers according to SCAV. Effective attack prompts that may be used in black-box attack scenarios can also be identified by using our optimization goal that balances ASR and response quality. Our comparative study, which involves human evaluations, confirms that our attack method significantly and consistently improves ASR (+8%~20%), language quality (+14%~42%), response usefulness (+8%~10%), and requires less training data.
@@
We further unveil the safety risks and mechanisms of diverse LLMs by applying our attack method. We discover that their safety issues are severe. For white-box LLMs, human evaluation results show that (a) even seemingly well-aligned LLMs like LLaMA-2 (Touvron et al., 2023) provide relevant answers to 96% to 100% malicious instructions; (b) the attacks of one LLM may be transferred to others, including black-box ones like GPT-4; and (c) existing methods aimed at making LLM unlearn harmful knowledge (Lu et al., 2024; Yao et al., 2023; Deng et al., 2023) may not have truly achieved their goals and still result in an ASR of 97.34% by attacking the unlearned LLM with our method.
@@ 2.1 Problem Formulation
Input: The attacks are generated by using an LLM $f$, whose embeddings are $\{\bm{e}^{1},...,\bm{e}^{L}\}$, where $\bm{e}^{l}\in\mathbb{R}^{d}$ is the embedding at the $l$-th layer, and $L$ is the number of layers. While we require model parameters of $f$ to interpret its safety mechanism and optimize attack performance on $f$, the generated attacks may also be applied to other LLMs or even black-box APIs, considering the potential attack transferability demonstrated in our experiments (Sections 3.3 and 4.3) and previous research (Liu et al., 2023; Zou et al., 2023b).
@@ 2.2 SCAV Framework
We first introduce our Safety Concept Activation Vector (SCAV) framework, which effectively guides the subsequent attack process by quantitatively interpreting LLMs’ embedding-level safety mechanisms. Specifically, given an embedding $\bm{e}$, we aim to estimate the probability $P_{\text{m}}({\bm{e}})$ that the LLM considers $\bm{e}$ as malicious [^1] We omit the superscript $l$ in $\bm{e}^{l}$ for conciseness when there is no ambiguity.. This is achieved by using Concept Activation Vector (Kim et al., 2018), a classic interpretation method that follows the linear interpretability assumption commonly used in existing interpretation methods (Alain and Bengio, 2016; Raghu et al., 2017; Bau et al., 2017; Smilkov et al., 2017). Specifically, it assumes that a deep model embedding $\bm{e}$ can be mapped to a concept that humans can understand (in our paper, the “safety” concept) after a linear transformation. Accordingly, the probability that the LLM considers $\bm{e}$ malicious can be modeled through a linear classifier:
@@
$\mathop{\arg\min}\limits_{\bm{w},b}-\frac{1}{|D|}\sum_{(y,\bm{e})\in D}[y\log P_{\text{m}}(\bm{e})+(1-y)\log(1-P_{\text{m}}(\bm{e})]+\frac{1}{2}||\bm{w}||^{2}+\frac{1}{2}b^{2}$ (2)
where $D$ is the training dataset, $y=1$ if the input instruction is malicious and is 00 if the instruction is safe. Like existing attack baselines that consider model inter workings (Zou et al., 2023a; Li et al., 2024), we also require a dataset with both malicious and safe instructions to determine the label $y$. However, we require much less training data (Figure 3), demonstrating the effectiveness of SCAV-based model interpretation that helps eliminate potentially misleading heuristics (Section 2.3.1).
@@
Verifying the linear interpretability assumption. To check whether the linear interpretability assumption holds for the safety concept in LLMs, we investigate the test accuracy of classifier $P_{\text{m}}$. A high accuracy means that the embeddings of malicious and safe instructions are linearly separatable in the LLM hidden space. As shown in Figure 1, for aligned LLMs (Vicuna and LLaMA), the test accuracy becomes larger than 95% starting from the 10th or 11th layer and grows to over 98% at the last layers. This indicates that a simple linear classifier can accurately interpret LLMs’ safety mechanism and that LLMs usually start to model the safety concept from the 10th or 11th layer. In contrast, the test accuracy of the unaligned LLM (Alpaca) is much lower.
@@ 2.3.1 Optimizing Attacks for a Single Layer
Given embedding $\bm{e}$ at an intermediate layer, we attack $\bm{e}$ by changing it to $\tilde{\bm{e}}=\bm{e}+\epsilon\cdot\bm{v}$, where $\epsilon\in\mathbb{R}$ is the perturbation magnitude and $\bm{v}\in\mathbb{R}^{d}$ ($||\bm{v}||=1$) is the perturbation direction. While existing white-box attack methods (Zou et al., 2023a; Li et al., 2024) heuristically determine the perturbation direction and provide no guidance for the perturbation magnitude, we optimize $\epsilon$ and $\bm{v}$ simultaneously by solving the following constrained optimization problem, which ensures small performance loss of LLMs and high attack success rate:
@@
The optimization problem in Equation (5) has a closed-form solution (proof in Appendix B):
@@
where ${\mathbb{I}}(\cdot)$ is an indicator function that transforms boolean values (false or true) into integers (0 or 1).
Method intuition and analysis of baselines. Our perturbation direction $\bm{v}$ is perpendicular to the hyperplane that separates malicious instructions from safe ones, according to Equation (4). As shown in Figure 2, this allows us to move the embeddings of malicious instructions to the subspace of safe instructions consistently with the shortest possible distance. In contrast, baselines RepE (Zou et al., 2023a) and JRE (Li et al., 2024) may result in ineffective perturbations. For example, the perturbation vector of JRE is perpendicular to the correct direction in Case 3, and RepE may generate opposite perturbations in different runs. This is caused by their potentially misleading heuristics. Both methods heuristically obtain a perturbation vector that depicts the global difference between embeddings of malicious instructions ($\bm{e}_{\text{m}}$) and embeddings of safe instructions ($\bm{e}_{\text{s}}$). This is achieved by randomly subtracting $\bm{e}_{\text{m}}$ and $\bm{e}_{\text{s}}$ and performing PCA analysis (Zou et al., 2023a) or dimension selection (Li et al., 2024) to identify a potentially interesting direction. Such a perturbation vector relies heavily on the global data distribution, requires more data points, and may not align with the hyperplane for separating $\bm{e}_{\text{m}}$ and $\bm{e}_{\text{s}}$, leading to attack failure (due to the large $P_{\text{m}}(\tilde{\bm{e}})$) or low-quality responses (due to perturbation in the wrong direction).
Figure 2: Comparison of perturbations added by our method (SCAV) and the baselines RepE (Zou et al., 2023a) and JRE (Li et al., 2024). Our method consistently moves embeddings of malicious instructions to the subspace of safe instructions, while the baselines may result in ineffective or even opposite perturbations.
@@ 2.4 Prompt-Level Attack
In this subsection, we demonstrate how our SCAV classifier $P_{\text{m}}$ can effectively guide the generation of an attack prompt $S$. Attack prompts can be combined with original user instructions to manipulate LLMs’ behavior. Existing white-box attack methods, such as GCG (Zou et al., 2023b) and AutoDAN (Liu et al., 2023), automatically generate adversarial prompts to maximize the probability of a certain target response $T$ (e.g., Sure, here is how to make a bomb). The heuristically determined target response $T$ is often different from the real positive response when an LLM is successfully attacked. There is no guarantee that the attack success rates can be accurately or completely estimated by using the generation probability of $T$, thereby limiting the performance of existing methods.
@@
where $\bm{e}^{L}$ is the last-layer embedding of a user instruction $x$, and $\bm{e}^{L}_{S}$ is the last-layer embedding when the attack prompt $S$ is combined with $x$ to manipulate the model. The first term $P_{\text{m}}(\bm{{e}}^{L}_{S})$ ensures the effectiveness of the attack, while the second term $||\bm{{e}}^{L}_{S}-\bm{e}^{L}||$ guarantees minimal modifications to the model in order to avoid low-quality model responses. We solve Equation (5) by using AutoDAN’s hierarchical genetic algorithm. We do not use the constrained formulation in Equation (3), because 1) it is not easy to incorporate constraints into the hierarchical genetic algorithm; and 2) it is difficult to determine $P_{0}$ here since we cannot directly control the embeddings to ensure a low value of $P_{\text{m}}$.
@@ 3.1 Experimental Setup
Baselines. We compare SCAV with four baselines: DeepInception (Li et al., 2023) uses manually-crafted attack prompts, AutoDAN (Liu et al., 2023) learns an attack prompt based on LLMs’ output logit distribution and its gradient, and RepE (Zou et al., 2023a) and JRE (Li et al., 2024) require model parameters and attack by changing LLM embeddings.
Datasets. The training data for our method, RepE, and JRE are 140 malicious instructions from Advbench (Chen et al., 2022) and HarmfulQA (Bhardwaj and Poria, 2023) and 140 safe instructions generated by utilizing GPT-4. Testing datasets are the subset version of Advbench (Chen et al., 2022) and StrongREJECT (Souly et al., 2024), which do not overlap with the training data, and each contains 50 malicious instructions covering diverse malicious scenarios.
Victim LLMs. We consider three well-aligned LLMs as our attacking targets: two white-box models LLaMA-2-7B/13B-Chat (Touvron et al., 2023) and one black-box API GPT-4-turbo-2024-04-09.
@@
The first category contains one commonly-used criterion ASR-keyword (Zou et al., 2023b), which computes the attack success rates by simple keyword matching. If any of the predefined refusal keywords (e.g., I cannot) appear in the output, the attack will be judged as failed, otherwise it will be judged as successful (see Appendix A for the keyword list). ASR-keyword may not always accurately reflect whether an attack is successful. For example, if an LLM generates garbled responses with no useful information due to too large a perturbation, ASR will still consider the attack successful, since no refusal keywords are present.
The second category of criteria is then designed, which leverages human judges or GPT4 to better evaluate response quality. Specifically, we consider three criteria: ASR-answer that evaluates whether LLMs give relevant answers to malicious instructions, ASR-useful that decides whether the responses are useful, and Language flaws that determines whether the responses contain language flaws such as repetition, inconsistency, or unspecific paragraphs. The complete definition of each criterion, prompt used for GPT-4, and user study information can be found in Appendix A and F.
Implementation details of our method and baselines are given in Appendix D.

### v5

@@ Abstract
Despite careful safety alignment, current large language models (LLMs) remain vulnerable to various attacks. To further unveil the safety risks of LLMs, we introduce a Safety Concept Activation Vector (SCAV) framework, which effectively guides the attacks by accurately interpreting LLMs’ safety mechanisms. We then develop an SCAV-guided attack method that can generate both attack prompts and embedding-level attacks with automatically selected perturbation hyperparameters. Both automatic and human evaluations demonstrate that our attack method significantly improves the attack success rate and response quality while requiring less training data. Additionally, we find that our generated attack prompts may be transferable to GPT-4, and the embedding-level attacks may also be transferred to other white-box LLMs whose parameters are known. Our experiments further uncover the safety risks present in current LLMs. For example, in our evaluation of seven open-source LLMs, we observe an average attack success rate of 99.14%, based on the classic keyword-matching criterion. Finally, we provide insights into the safety mechanism of LLMs. The code is available at https://github.com/SproutNan/AI-Safety_SCAV.
@@ 1 Introduction
The rapid advancement in large language models (LLMs) has raised significant concerns about their potential misuse  [1; 2; 3; 4]. Developers usually conduct intensive alignment work [5; 6; 7; 8; 9; 10; 11; 12] to prevent powerful LLMs from being exploited for harmful activities. However, researchers have discovered that these time-consuming safety alignment efforts can be reversed by various attack methods [13; 14; 15; 16]. These methods can identify vulnerabilities in safety alignment technologies and enable developers to fix them promptly, reducing the societal safety risks of LLMs.
@@
Existing attack methods utilize different levels of information from LLMs to achieve varying degrees of model understanding and control. Pioneering attack methods manually design prompt templates [15; 17] or learn attack prompts without information about intermediate layers of LLMs [13; 14]. The attack prompts may be applied to various LLMs, supporting both black-box attacks on APIs and white-box scenarios where model parameters are released. However, their attack success rates (ASR) [14] are constrained by an insufficient understanding of LLMs’ internal working mechanisms. Some recent attack works further utilize model embeddings at intermediate layers [16; 18]. By better understanding models’ safety mechanisms and perturbing relevant dimensions in the embeddings, these methods achieve significantly higher ASR on white-box LLMs. However, they cannot be applied to black-box APIs. Moreover, existing methods perturb LLM embeddings based on potentially misleading heuristics (Section 2.3.1). Due to the lack of a principled optimization goal, they result in a suboptimal ASR, may generate low-quality (e.g., repetitive) text, and require time-consuming grid search to find a good combination of hyperparameters (e.g., perturbation magnitude and layers).
@@
We aim to solve the aforementioned issues by more accurately modeling LLMs’ safety mechanisms, based on which principled optimization goals can be developed to well balance ASR and response quality, enable automatic hyperparameter selection, and support both embedding-level and prompt-level attacks. Specifically, we make the following contributions.
@@
First, we establish a Safety Concept Activation Vector (SCAV) framework that effectively guides the subsequent attack process by accurately interpreting LLMs’ safety mechanisms. It quantifies the probability that an LLM considers an embedding as malicious based on the concept activation vector [19], which linearly separates embeddings of malicious and safe instructions.
@@
We then develop an SCAV-guided attack method, which utilizes SCAV to design principled optimization goals for finding high-quality embedding-level and prompt-level attacks. Our embedding-level attack method eliminates the time-consuming hyperparameter tuning process by providing a closed-form solution of the optimal perturbation magnitudes and automatically selecting layers according to SCAV. Effective attack prompts that may be used in black-box attack scenarios can also be identified by using our optimization goal that balances ASR and response quality. Our comparative study, which involves human evaluations, confirms that our attack method significantly and consistently increases ASR (+8%~20%), language quality (+14%~42%), response usefulness (+8%~10%), and requires less training data.
@@
We further unveil the safety risks and mechanisms of diverse LLMs by applying our attack method. We discover that their safety issues are severe. For white-box LLMs, human evaluation results show that (a) even seemingly well-aligned LLMs like LLaMA-2 [20] provide relevant answers to 96% to 100% malicious instructions; (b) the attacks of one LLM may be transferred to others, including black-box ones like GPT-4; and (c) existing methods aimed at making LLM unlearn harmful knowledge [21; 22; 23] may not have truly achieved their goals and still result in an ASR of 97.34% by attacking the unlearned LLM with our method.
@@ 2.1 Problem Formulation
Input: The attacks are generated by using an LLM $f$, whose embeddings are $\{\bm{e}^{1},...,\bm{e}^{L}\}$, where $\bm{e}^{l}\in\mathbb{R}^{d}$ is the embedding at the $l$-th layer, and $L$ is the number of layers. While we require model parameters of $f$ to interpret its safety mechanisms and optimize attack performance on $f$, the generated attacks may also be applied to other LLMs or even black-box APIs, considering the potential attack transferability demonstrated in our experiments (Sections 3.3 and 4.3) and previous research [13; 14].
@@ 2.2 SCAV Framework
We first introduce our Safety Concept Activation Vector (SCAV) framework, which effectively guides the subsequent attack process by quantitatively interpreting LLMs’ embedding-level safety mechanisms. Specifically, given an embedding $\bm{e}$, we aim to estimate the probability $P_{\text{m}}({\bm{e}})$ that the LLM considers $\bm{e}$ as malicious [^1] We omit the superscript $l$ in $\bm{e}^{l}$ for conciseness when there is no ambiguity.. This is achieved by using Concept Activation Vector [19], a classic interpretation method that follows the linear interpretability assumption commonly used in existing interpretation methods [24; 25; 26; 27; 28; 29; 30; 31]. Specifically, it assumes that a deep model embedding $\bm{e}$ can be mapped to a concept that humans can understand (in our paper, the “safety” concept) after a linear transformation. Accordingly, the probability that the LLM considers $\bm{e}$ malicious can be modeled through a linear classifier:
@@
$\mathop{\arg\min}\limits_{\bm{w},b}-\frac{1}{|D|}\sum_{(y,\bm{e})\in D}[y\log P_{\text{m}}(\bm{e})+(1-y)\log(1-P_{\text{m}}(\bm{e})]$ (2)
where $D$ is the training dataset, $y=1$ if the input instruction is malicious and is 00 if the instruction is safe. Implementation details can be found at Appendix E.1. Like existing attack baselines that consider model inter workings [16; 18], we also require a dataset with both malicious and safe instructions to determine the label $y$. However, we require much less training data (Figure 3), demonstrating the effectiveness of SCAV-based model interpretation that helps eliminate potentially misleading heuristics (Section 2.3.1).
@@
Verifying the linear interpretability assumption. To check whether the linear interpretability assumption holds for the safety concept in LLMs, we investigate the test accuracy of classifier $P_{\text{m}}$. A high accuracy means that the embeddings of malicious and safe instructions are linearly separatable in the LLM hidden space. As shown in Figure 1, for aligned LLMs (Vicuna and LLaMA-2), the test accuracy becomes larger than 95% starting from the 10th or 11th layer and grows to over 98% at the last layers. This indicates that a simple linear classifier can accurately interpret LLMs’ safety mechanism and that LLMs usually start to model the safety concept from the 10th or 11th layer. In contrast, the test accuracy of the unaligned LLM (Alpaca) is much lower. We provide similar results on other LLMs in Appendix D.1.
@@ 2.3.1 Optimizing Attacks for a Single Layer
Given embedding $\bm{e}$ at an intermediate layer, we attack $\bm{e}$ by changing it to $\tilde{\bm{e}}=\bm{e}+\epsilon\cdot\bm{v}$, where $\epsilon\in\mathbb{R}$ is the perturbation magnitude and $\bm{v}\in\mathbb{R}^{d}$ ($||\bm{v}||=1$) is the perturbation direction. While existing white-box attack methods [16; 18] heuristically determine the perturbation direction and provide no guidance for the perturbation magnitude, we optimize $\epsilon$ and $\bm{v}$ simultaneously by solving the following constrained optimization problem, which ensures small performance loss of LLMs and high attack success rates:
@@
The optimization problem in Equation (5) has a closed-form solution (proof in Appendix C):
@@
where ${\mathbb{I}}(\cdot)$ is an indicator function that transforms false or true into 0 or 1.
Method Intuition and Analysis of Baselines. Our perturbation direction $\bm{v}$ is perpendicular to the hyperplane that separates malicious instructions from safe ones, according to Equation (4). As shown in Figure 2, this allows us to move the embeddings of malicious instructions to the subspace of safe instructions consistently with the shortest possible distance. In contrast, baselines RepE [16] and JRE [18] may result in ineffective perturbations. For example, the perturbation vector of JRE is perpendicular to the correct direction in Case 3, and RepE may generate opposite perturbations in different runs. This is caused by their potentially misleading heuristics. Both methods heuristically obtain a perturbation vector that depicts the global difference between embeddings of malicious instructions ($\bm{e}_{\text{m}}$) and embeddings of safe instructions ($\bm{e}_{\text{s}}$). This is achieved by randomly subtracting $\bm{e}_{\text{m}}$ and $\bm{e}_{\text{s}}$ and performing PCA analysis [16] or dimension selection [18] to identify a potentially interesting direction. Such a perturbation vector relies heavily on the global data distribution, requires more data points, and may not align with the hyperplane for separating $\bm{e}_{\text{m}}$ and $\bm{e}_{\text{s}}$, leading to attack failure (due to the large $P_{\text{m}}(\tilde{\bm{e}})$) or low-quality responses (due to perturbation in the wrong direction).
Figure 2: Comparison of perturbations added by our method (SCAV) and the baselines RepE [16] and JRE [18]. Our method consistently moves embeddings of malicious instructions to the subspace of safe instructions, while the baselines may result in ineffective or even opposite perturbations.
@@ 2.4 Prompt-Level Attack
In this subsection, we demonstrate how our SCAV classifier $P_{\text{m}}$ can effectively guide the generation of an attack prompt $S$. Attack prompts can be combined with original user instructions to manipulate LLMs’ behavior. Existing white-box attack methods, such as GCG [14] and AutoDAN [13], automatically generate adversarial prompts to maximize the probability of a certain target response $T$ (e.g., Sure, here is how to make a bomb). The heuristically determined target response is often different from the real positive response when an LLM is successfully attacked. There is no guarantee that the attack success rates can be accurately or completely estimated by using the generation probability of $T$, thereby limiting the performance of existing methods.
@@
where $\bm{e}^{L}$ is the last-layer embedding of a user instruction $x$, and $\bm{e}^{L}_{S}$ is the last-layer embedding when the attack prompt $S$ is combined with $x$ to manipulate the model. The first term $P_{\text{m}}(\bm{{e}}^{L}_{S})$ ensures the effectiveness of the attack, while the second term $||\bm{{e}}^{L}_{S}-\bm{e}^{L}||$ guarantees minimal modifications to the model in order to avoid low-quality model responses. We solve Equation (5) by using AutoDAN’s hierarchical genetic algorithm (See Appendix E.2.1 for details). We do not use the constrained formulation in Equation (3), because 1) it is not easy to incorporate constraints into the hierarchical genetic algorithm; and 2) it is difficult to determine $P_{0}$ here since we cannot directly control the embeddings to ensure a low value of $P_{\text{m}}$. See Appendix E.2.2 for more discussions of the design choice.
@@ 3.1 Experimental Setup
Baselines. We compare SCAV with the following baselines, which involve different kinds of LLM attacking paradigms.
  • DeepInception [17], which uses manually-crafted attack prompts.
  • AutoDAN [13] and GCG [14], which learn attack prompts based on LLMs’ output logit distribution and gradient.
  • RepE [16] and JRE [18], which require model parameters and attack by changing LLM embeddings.
  • Soft prompt [32], which also enables attacking LLMs in embedding space.
Datasets. The training data for embedding-level attacks are 140 malicious instructions from Advbench [33] and HarmfulQA [34] and 140 safe instructions generated by utilizing GPT-4. Testing datasets are the subset version of Advbench [33] and StrongREJECT [35], which do not overlap with the training data, and each contains 50 malicious instructions covering diverse malicious scenarios.
Victim LLMs. We consider three well-aligned LLMs as our attacking targets: two white-box models LLaMA-2-7B/13B-Chat [20] and one black-box API GPT-4-turbo-2024-04-09.
@@
  • ASR-keyword [14], which is a commonly used criterion for computing attack success rate (ASR) via simple keyword matching. If any of the predefined refusal keywords (e.g., I cannot) appears in the output, the attack will be judged as failed, otherwise it will be judged as successful (see Appendix B for the keyword list). ASR-keyword may not always accurately reflect whether an attack is successful. For example, if an LLM generates garbled responses with no useful information due to a large perturbation, ASR will still consider the attack successful, since no refusal keywords are present.
  • LLM-based Criteria, including three proposed criteria: (1) ASR-answer that evaluates whether LLMs give relevant answers to malicious instructions, (2) ASR-useful that decides whether the responses are useful, and (3) Language flaws that determines whether the responses contain language flaws such as repetition, inconsistency, or unspecific paragraphs. In general, ASR-useful is a more strict criterion than ASR-answer in harmfulness evaluation. These newly designed criteria leverage human judges or GPT-4 to better evaluate response quality.
The complete definition of each criterion, prompt used for GPT-4, and user study information can be found in Appendix B and H. The implementation details of our method, baselines and comparative experiments are given in Appendix E.